### v2

# Conditional Denoising Diffusion Probabilistic Models for Data Reconstruction Enhancement in Wireless Communications

Abstract
In this paper, conditional denoising diffusion probabilistic models (DDPMs) are proposed to enhance the data transmission and reconstruction over wireless channels. The underlying mechanism of DDPM is to decompose the data generation process over the so-called “denoising” steps. Inspired by this, the key idea is to leverage the generative prior of diffusion models in learning a “noisy-to-clean” transformation of the information signal to help enhance data reconstruction. The proposed scheme could be beneficial for communication scenarios in which a prior knowledge of the information content is available, e.g., in multimedia transmission. Hence, instead of employing complicated channel codes that reduce the information rate, one can exploit diffusion priors for reliable data reconstruction, especially under extreme channel conditions due to low signal-to-noise ratio (SNR), or hardware-impaired communications. The proposed DDPM-assisted receiver is tailored for the scenario of wireless image transmission using MNIST dataset. Our numerical results highlight the reconstruction performance of our scheme compared to the conventional digital communication, as well as the deep neural network (DNN)-based benchmark. It is also shown that more than 10 dB improvement in the reconstruction could be achieved in low SNR regimes, without the need to reduce the information rate for error correction.
Index Terms:
AI-native wireless, diffusion priors, generative AI, network resilience, wireless AI, conditional denoising diffusion models.
I Introduction
I-A Motivation & Background
With the emergence of deep generative models, the realm of artificial intelligence (AI) has witnessed a paradigm shift towards the novel concept of generative AI that can facilitate the development of AI-based systems [2]. At the same time, from data communication and networking perspective, “connected intelligence” is envisioned as the most significant driving force in the sixth generation (6G) of communications—AI and machine learning (AI/ML) algorithms are envisioned to be widely incorporated into 6G wireless networks, realizing “AI-native” communication systems [3, 5, 4, 6]. This underscores the need for novel AI/ML-based solutions to be tailored to meet the requirements of emerging communication scenarios.
The evolution of diffusion models [7], as the new paradigm in deep generative models, can be regarded as one of the main factors contributing to the recent breakthroughs in generative AI, that have already showcased notable success with some of the popular examples such as DALL.E 2 by OpenAI, [^1] https://openai.com/dall-e-2 and ImageGen by Google Brain. [^2] https://imagen.research.google/ Diffusion models have realized unprecedented results in different applications, such as computer vision, natural language processing (NLP), data analysis and synthesis, and medical imaging [8]. Please see [9] and references therein for a comprehensive survey on diffusion models. The core concept of diffusion modelling is that if we could develop an ML model that is capable of learning the systematic decay of information, due to noise/distortions, then it should be possible to “reverse” the process and recover the information back from the noisy/erroneous data. This is fundamentally different from conventional generative models, like generative adversarial networks (GANs) and variational autoencoders (VAEs). Accordingly, the close underlying relation between the key concepts on how diffusion models work and the problems in wireless communication systems has motivated us to carry out this research.
I-B Related Works
Ongoing research on diffusion models encompasses both theoretical advancements and practical applications across different domains of computer science [9]. However, there have been only a few papers in wireless communication literature that have started looking into the potential merits of diffusion-based generative models for wireless systems [10, 11, 12, 13, 14, 15]. The authors in [10] study a workflow for utilizing diffusion models in wireless network management. They study the incentive mechanisms for generating contracts using diffusion models in mobile AI-generated content services. They exploit the exploration capability of diffusion models in Q-learning algorithms for network management. Denoising diffusion model is utilized in [11] to generate synthetic channel realizations conditioned on the message information. The authors tackle the problem of differentiable channel model within the gradient-based training process of end-to-end ML-based communications. The results highlight the performance of diffusion models as an alternative to generative adversarial network (GAN)-based schemes. It is shown in [11] that GANs experience unstable training and less diversity in generation performance, due to their adversarial training nature. However, diffusion models maintain a more stable training process and a better generalization during inference. Noise-conditioned score network (NCSN)-based channel estimation is studied in [12] and [13] for multi-input-multi-output (MIMO) wireless communication scenario. The authors estimate the gradient of the high-dimensional log-prior of wireless channels, using RefineNet neural network architecture. Posterior sampling method is further applied to make the generated channel estimations consistent with the pilot signals measured at the receiver. Diffusion models have also been incorporated into the so-called semantic communication systems in [15, 14]. More specifically, deep learning-based joint source-channel coding (Deep-JSCC) is combined with diffusion models to complement semantic data transmission schemes with a generative component. The initial results show that the perceptual quality of reconstruction can be improved in such scenarios. Nevertheless, those schemes in [15, 14] propose to employ two different types of deep learning models each maintaining a distinct objective function (while their optimization also depends on each other), i.e., the autoencoder for semantic transmission and reception, and the diffusion models for reconstruction enhancement. This can impose serious concerns in terms of the computational overhead to the network. In addition, the output signal of a neural encoder (i.e., the channel input) does not necessarily follow the standard format of modulation signals, which makes it challenging to employ semantic transmission ideas in a real-world communication scenario.
I-C Our Contributions
In this paper, we propose conditional denoising diffusion probabilistic models (DDPMs) for enhancing the wireless image transmission in digital communication schemes. Different from GANs and VAEs, the underlying mechanism of DDPM is to first transform the ground-truth signal into an isotropic Gaussian distribution in the forward diffusion process, and then try to regenerate data over the so-called “denoising” steps by removing the noise introduced in the diffusion process. DDPMs in their vanilla format are based upon the assumption that the denoising process starts with isotropic Gaussian noise [7]. However, in a practical wireless system, the received/decoded signals are a degraded version of the ground-truth information signals, not necessarily a pure isotropic Gaussian noise, while the degradation is different in different distortion conditions such as different SNRs and transceiver specifications. To address this, we employ conditional diffusion probabilistic models, in which the noisy degraded signals are incorporated into the denoising framework of the diffusion model. This way, the diffusion model learns to reconstruct the ground-truth information from the degraded received signal, instead of a pure noise.
In our scheme, the generative prior of conditioned DDPMs is exploited to enhance data reconstruction at the receiver of a communication system. This is particularly pronounced in extreme channel conditions with poor connectivity due to low signal-to-noise ratio (SNR), or hardware-impaired communications. In such scenarios, error correcting codes might not be able to correct the mismatches, or complicated codes might be needed at the cost of decreasing the information rate to correct the errors. Instead, the proposed diffusion-aided scheme can be utilized for reliable data reconstruction without the need for decreasing the information rate. This can be tailored for the communication scenarios, in which a prior knowledge of the information signals is available that can be exploited by a DDPM to learn the “denoising” process.
The proposed DDPM-assisted receiver is employed in conjunction with the NVIDIA Sionna simulator [20], by plugging the trained diffusion model to the receiver side, highlighting the compatibility of the scheme with a practical communication system. Moreover, in contrast to [15, 14], the proposed scheme does not rely on any additional autoencoder model. Numerical results highlight that the resilience of the wireless system can be improved against practical non-idealities such as hardware impairments and channel distortions. We compare the reconstruction performance of our scheme to the conventional digital communications without diffusions, and also the deep neural network (DNN)-based benchmark [18]. It is also shown that more than 10 dB improvement in the reconstruction could be achieved in low SNR regimes, without the need to reduce the information rate for error correction.
I-D Paper Organization and Notations
The rest of the paper is organized as follows. In Section II, we first introduce the concept of DDPMs, the forward, and the reverse diffusion processes, together with the main formulas and the corresponding loss functions. We also try to point out insightful intuitions behind the main formulas of the DDPM framework. Our system model is introduced in Section III. We first provide the generic formulation of a communication system under non-linear HWIs. We then address the proposed conditional DDPM by modifying the vanilla diffusion model. Numerical experiments are provided in Section IV. Finally, Section V concludes the paper.
Notations
Vectors and matrices are represented, respectively, by bold lower-case and upper-case symbols. Absolute value of a scalar variable and the $\ell_{2}$ norm of a vector are denoted, respectively by ||$\cdot$|| ⋅ | and |||||$\cdot$||| | ⋅ | |. Notation $\mathcal{N}(\bm{x};\bm{\mu},\bm{\Sigma})$ stands for the multivariate normal distribution with mean vector $\bm{\mu}$ and covariance matrix $\bm{\Sigma}$ for a random vector $\bm{x}$. Similarly, complex normal distribution with the corresponding mean vector and covariance matrix is denoted by $\mathcal{CN}(\bm{\mu},\bm{\Sigma})$. Moreover, the expected value of a random variable is denoted by $\mathbb{E}\left[\cdot\right]$ Sets are denoted by calligraphic symbols, and $\bm{0}$ and $\bf I$ show all-zero vector and identity matrix of the corresponding size, respectively. Moreover, $[N]$, (with $N$ as integer) denotes the set of all integer values from $1$ to $N$, and $\mathsf{Unif}[N]$, $N>1$, denotes discrete uniform distribution with samples between $1$ to $N$. Real and imaginary parts of a complex-valued vector are denoted by ${}\Re\{\cdot\}$ and ${}\Im\{\cdot\}$, respectively.
Figure 1: Diffusion models: A general overview. A Markov chain is defined to mimic the forward diffusion steps, during which random perturbation noise is purposefully added to the original data. Then in a reverse process, the model learns to construct the desired data samples out of noise.
II Preliminaries on Denoising Diffusion Probabilistic Models
Diffusion models are a new family of state-of-the-art probabilistic generative models which are inspired by non-equilibrium thermodynamics [16]. They are comprised of two processes, namely the forward diffusion process and the parametric reverse diffusion process, as illustrated in Fig. 1. During the forward process, samples are diffused using Gaussian kernels to add noise until the fully-distorted signals follow an isotropic Gaussian distribution. Then in the reverse process, the model tries to “decode” the data via “denoising” the perturbed samples in a hierarchical fashion. In this paper, we focus on DDPMs a the new state-of-the-art family of diffusion models. In what follows, we study the concept of DDPMs together with the main formulas and the corresponding loss functions. We also try to point out insightful intuitions behind the underlying formulas of the DDPM framework.
II-A Forward Diffusion Process
Let ${\bm{x}}_{0}$ be a data sample from some distribution ${q}({\bm{x}}_{0})$, i.e., ${\bm{x}}_{0}\sim{q}({\bm{x}}_{0})$. For a finite number, $T$, of time-steps, the forward diffusion process $q({\bm{x}}_{t}|{\bm{x}}_{t-1})$ is defined by adding Gaussian noise at each time-step $t\in[T]$ according to a known “variance schedule” $10<\beta_{1}<\beta_{2}<\cdots<\beta_{T}<$ 1. This can be formulated as
$q(\bm{x}_{t}|\bm{x}_{t-1})$ $\sim\mathcal{N}(\bm{x}_{t};\sqrt{1-\beta_{t}}\bm{x}_{t-1},\beta_{t}\bm{I}),$ (1)
$q(\bm{x}_{1:T}|\bm{x}_{0})$ $=\prod^{T}_{t=1}q(\bm{x}_{t}|\bm{x}_{t-1}).$ (2)
Invoking (2), the data sample gradually loses its distinguishable features as the time-step goes on, where with $T\!\rightarrow\!\infty$, ${\bm{x}}_{T}$ approaches an isotropic Gaussian distribution with covariance matrix ${\bf\Sigma}\!=\!\sigma^{2}\bm{I}$ for some $\sigma\!>\!0$ [7]. According to (1), each new sample at time-step $t$ can be drawn from a conditional Gaussian distribution with mean vector ${\bm{\mu}}_{t}=\sqrt{1-\beta_{t}}\bm{x}_{t-1}$ and covariance matrix ${\bf\Sigma}^{2}_{t}=\beta_{t}\bf I$. Hence, the forward process is realized by sampling from a Gaussian noise $\bm{\epsilon}_{t-1}\sim{\cal N}(\bf 0,I)$ and setting
${\bm{x}}_{t}=\sqrt{1-\beta_{t}}$ ${\bm{x}}_{t-1}+\sqrt{\beta_{t}}{\bm{\epsilon}}_{t-1}.$ (3)
A useful property for the forward process in (3) is that we can sample ${\bm{x}}_{t}$ at any arbitrary time-step $t$ in a closed-form expression, through recursively applying the reparameterization trick from ML literature [19], and considering the sum properties of two Gaussian random variables. This results in the following formulation
$\bm{x}_{t}$ $=\sqrt{\bar{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{0},$ (4)
$q({\bm{x}}_{t}|{\bm{x}}_{0})$ $\sim\mathcal{N}\left({\bm{x}}_{t};\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0},(1-\bar{\alpha}_{t})\bm{I}\right),$ (5)
$\bar{\alpha}_{t}\!$ $=\!\prod_{i=1}^{t}(1-\alpha_{i}),\quad\alpha_{t}=1-\beta_{t}.$ (6)
Notably, (4) implies that one can directly sample ${\bm{x}}_{t}$, $t\in[T]$ from $x_{0}$.
II-B Parametric Reverse Diffusion Process
Now the problem is to reverse the process in (4) and sample from $q(\bm{x}_{t-1}|\bm{x}_{t})$, so that we regenerate the true samples from some Gaussian noise $\bm{x}_{T}$. According to [16], for $\beta_{t}$ small enough, $q(\bm{x}_{t-1}|\bm{x}_{t}),\forall t\in[T]$ also follows Gaussian distribution. However, we cannot easily estimate the distribution, since it requires knowing the distribution of all possible data samples (or equivalently exploiting the entire dataset) to be able to calculate the conditional probability. Hence, to approximate the conditional probabilities and run the reverse diffusion process, we need to learn a probabilistic model $p_{\bm{\theta}}(\bm{x}_{t-1}|\bm{x}_{t})$ that is parameterized by ${\bm{\theta}}$. Accordingly, the following expressions can be written
$p_{\bm{\theta}}(\bm{x}_{t-1}|\bm{x}_{t})$ $\sim\mathcal{N}(\bm{x}_{t-1};\bm{\mu}_{\bm{\theta}}(\bm{x}_{t},t),\bm{\Sigma}_{\bm{\theta}}(\bm{x}_{t},t)),$ (7)
$p_{\bm{\theta}}(\bm{x}_{0:T})$ $=p(\bm{x}_{T})\prod^{T}_{t=1}p_{\bm{\theta}}(\bm{x}_{t-1}|\bm{x}_{t}).$ (8)
As can be seen from (7), the mean and variance of the reverse process are also conditioned on the time-step $t$. Now the problem simplifies to learning the mean vector $\bm{\mu}_{\bm{\theta}}(x_{t},t)$ and the covariance matrix $\bm{\Sigma}_{\bm{\theta}}(x_{t},t)$ for the probabilistic model $p_{\bm{\theta}}(\cdot)$, where a neural network can be trained to approximate (learn) the reverse process with ${\bm{\theta}}$ showing the parameters of the neural network.
Before proceeding with the details of learning the reverse diffusion process, we note that if we condition the reverse process on ${\bm{x}}_{0}$, this conditional probability becomes tractable. The intuition behind this could be explained as follows. A painter (our generative model) requires a reference image ${\bm{x}}_{0}$ to be able to gradually draw a picture. Hence, when we have ${\bm{x}}_{0}$ as a reference, we can take a small step backwards from noise to generate the data samples, and the reverse step (drawing a picture in our metaphor) would be formulated as $q({\bm{x}}_{t-1}|{\bm{x}}_{t},{\bm{x}}_{0})$. Mathematically speaking, by knowing the conditional probabilities of $q({\bm{x}}_{t}|{\bm{x}}_{0})$ and $q({\bm{x}}_{t-1}|{\bm{x}}_{0})$, and utilizing Bayes rule, we can derive $q({\bm{x}}_{t-1}|{\bm{x}}_{t},{\bm{x}}_{0})$ with a similar expression to (7)
$q(\bm{x}_{t-1}|\bm{x}_{t},\bm{x}_{0})$ $\sim\mathcal{N}(\bm{x}_{t-1};\hskip 4.2679pt{\tilde{\bm{\mu}}}(\bm{x}_{t},\bm{x}_{0},t),{\tilde{\beta}_{t}}\bm{I}),$ (9)
${\tilde{\bm{\mu}}}(\bm{x}_{t},\bm{x}_{0},t)$ $=\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}{\bm{x}}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}{\bm{x}}_{0},$ (10)
${\tilde{\beta}_{t}}$ $=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}.$ (11)
Invoking (9)–(11), one can infer that the covariance matrix in (9) has no learnable parameter. Hence, we simply need to learn the mean vector ${\tilde{\bm{\mu}}}(\bm{x}_{t},\bm{x}_{0},t)$. In the following, we first simplify (10), and then propose the DDPM-based training framework.
Thanks to the reparameterization trick and with a similar approach to (4), we can express ${\bm{x}}_{0}$ as follows
$\bm{x}_{0}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}(\bm{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{t}).$ (12)
Substituting ${\bm{x}}_{0}$ in (10) by (12) results in
[…]
Now we can learn the conditioned probability distribution $p_{\bm{\theta}}(\bm{x}_{t-1}|\bm{x}_{t})$ of the reverse diffusion process by training a neural network that approximates $\tilde{\bm{\mu}}(\bm{x}_{t},\bm{x}_{0},t)$. Therefore, we simply need to set the approximated mean vector $\bm{\mu}_{\bm{\theta}}(\bm{x}_{t},t)$ to have the same form as the target mean vector $\tilde{\bm{\mu}}(\bm{x}_{t},\bm{x}_{0},t)$. Since $\bm{x}_{t}$ is available at time-step $t$, we can reparameterize the neural network to make it approximate $\bm{\epsilon}_{t}$ from the input $\bm{x}_{t}$. Compiling these facts results in the following expression for $\bm{\mu}_{\bm{\theta}}(\bm{x}_{t},t)$
$\bm{\mu}_{\bm{\theta}}(\bm{x}_{t},t)$ $={\frac{1}{\sqrt{\alpha_{t}}}\Big{(}\bm{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{t},t)\Big{)}},$ (14)
where $\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{t},t)$ denotes the model to predict $\bm{\epsilon}_{t}$.
We now define the loss function $\mathcal{L}_{t}$ for time-step $t$, aiming to minimize the difference between $\bm{\mu}_{\bm{\theta}}(\bm{x}_{t},t)$ and $\tilde{\bm{\mu}}(\bm{x}_{t},\bm{x}_{0},t)$.
$\mathcal{L}_{t}$ $={\mathbb{E}}_{\begin{subarray}{l}t\sim{\mathsf{Unif}}[T]\\ {\bm{x}}_{0}\sim q({\bm{x}}_{0})\\ \bm{\epsilon}_{0}\sim\mathcal{N}(0,\bm{I})\\ \end{subarray}}\Big{[}\|\bm{\epsilon}_{t}-\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{t},t)\|^{2}\Big{]}$
$={\mathbb{E}}_{\begin{subarray}{l}t\sim{\mathsf{Unif}}[T]\\ {\bm{x}}_{0}\sim q({\bm{x}}_{0})\\ \bm{\epsilon}_{0}\sim\mathcal{N}(0,\bm{I})\\ \end{subarray}}\Big{[}\|\bm{\epsilon}_{t}-\bm{\epsilon}_{\bm{\theta}}(\sqrt{\bar{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{t},t)\|^{2}\Big{]}.$ (15)
Invoking (14) and (15), at each time-step $t$, the DDPM model takes $\bm{x}_{t}$ as input and returns the distortion components $\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{t},t)$. Moreover, $\bm{\epsilon}_{t}$ denotes the diffused noise term at time-step $t$.
III System Model and Proposed Scheme
III-A System Model & Formulas
Consider a point-to-point communication system, where a source node (the transmitter) aims to transmit the ground-truth information $x_{0}\sim q(x_{0})$ to a destination node (the receiver) over the communication channel. We denote by $s$, the channel input signal subject to the power constraint $\mathbb{E}[|s|^{2}]\leq P$, for which we can write $s=g(x_{0})$ with $g(\cdot)$ denoting the mapping from information signals to the channel input (including the off-the-shelf modulation and coding functionalities). The information-bearing signal $s$ is then sent over the communication channel and $y=\eta(s)$ is observed at the receiver, where $\eta(\cdot)$ stands for the channel distortion and noise, as well as the impairments caused by the non-ideal transceiver hardware. To model $\eta(\cdot)$, we consider a practical scenario of hardware-impaired communication, where we employ the well-known generic communication model of [21]. Denoting the Rayleigh fading wireless channel coefficient by $h$, we have
${y}$ $={h}(\sqrt{P}{s}+{\tau}^{t})+\tau^{r}+n,$ (16)
$\tau^{t}$ $\sim\mathcal{CN}({0},\kappa^{t}P),\quad\tau^{r}\sim\mathcal{CN}(0,\kappa^{r}P|h|^{2}),$ (17)
where $P$ denotes the transmit power, $\tau^{t}$ is the distortion noise caused by the transmitter hardware with the corresponding impairment level $\kappa^{t}$, and $\tau^{r}$ reflects the hardware distortion at the receiver with $\kappa^{r}$ showing the level of impairment at the receiver hardware. Notably, ${\tau^{r}}$ is conditionally Gaussian, given the channel realization ${h}$. This is a well-known and experimentally-validated model for HIs, which is widely-adopted in wireless communication literature [21]. We emphasize that according to [21], the distortion noise caused at each radio frequency (RF) device is proportional to its signal power, where a fixed portion of the information signal is turned into distortion noise due to quantization errors in analog-to-digital converter (ADC), inter-carrier interference induced by phase noise, leakage from the mirror subcarrier under I/Q imbalance, nonlinearities in power amplifiers, etc. To be aligned with the “batch-processing” nature of AI/ML algorithms and wireless communications simulators [20], we reformulate the signaling expressions in matrix-based format. To do so, we define the batch (with size $K$) of information-bearing samples as ${\bm{s}}\overset{\Delta}{=}\left[s_{1},s_{2},\cdots,s_{K}\right]^{\sf T}$ corresponding to the batch of ground-truth information samples ${\bm{x}_{0}}\overset{\Delta}{=}\left[x_{0}^{(1)},x_{0}^{(2)},\cdots,x_{0}^{(K)}\right]^{\sf T}$ with the underlying distribution $\bm{x}_{0}\sim q(\bm{x}_{0})$. Hence, the batch of received signals at the receiver can be expressed as
$\bm{y}=\sqrt{P}\hskip 2.84526pt\bm{H}\bm{s}+\bm{\zeta},$ (18)
where $\bm{H}=\hskip 2.84526pt\mathsf{diag}(\bm{h})$, ${\bm{h}}\overset{\Delta}{=}\left[h^{(1)},h^{(2)},\ldots,h^{(K)}\right]^{\sf T}$ stands for the channel realizations vector corresponding to the transmission of information samples, and ${\bm{\zeta}}\overset{\Delta}{=}\bm{H}\bm{\tau}^{t}+\bm{\tau}^{r}+\bm{n}$ represents the effective noise-plus-distortion at the receiver with the conditional distribution $\bm{\zeta}|{\bm{h}}\sim\mathcal{CN}({{\bf 0}}_{K},{\bm{\Sigma}})$ and the covariance matrix ${\bf\Sigma}$ given by ${\bf\Sigma}=P(\kappa^{t}+\kappa^{r})\bm{G}+\sigma^{2}\bm{I}_{K}$. We also define
${\bm{\tau}}^{t}$ =$\overset{\Delta}{=}\left[\tau^{t}_{1},\tau^{t}_{2},\ldots,\tau^{t}_{K}\right]^{\sf T}\sim\mathcal{CN}({{\bf{0}}_{K}},\kappa^{t}P{\bm{I}}_{K}),$ (19)
${\bm{\tau}}^{r}$ =$\overset{\Delta}{=}\left[\tau^{r}_{1},\tau^{r}_{2},\ldots,\tau^{r}_{K}\right]^{\sf T}\sim\mathcal{CN}({\bf 0}_{K},\kappa^{r}P{\bm{G}}),$ (20)
which is conditionally Gaussian given the channel realizations $\{{h}_{k}\}_{k\in[K]}$, with $\bm{G}=\mathsf{diag}(\left[|h_{1}|^{2},|h_{2}|^{2},\ldots,|h_{K}|^{2}\right]^{\sf T})$. Due to the fact that neural networks can only process real-valued inputs, we map complex-valued symbols to real-valued tensors, and rewrite (18) by stacking the real and imaginary components. This results in the following expression.
$\tilde{\bm{y}}=\tilde{\bm{H}}\tilde{\bm{s}}+\bm{\nu},$ (21)
where
$\tilde{\bm{y}}$ =[$=\begin{bmatrix}\Re\{{\bm{y}}\}\\ \Im\{{\bm{y}}\}\end{bmatrix}\in\mathbb{R}^{2K},$ (22)
$\tilde{{\bm{s}}}$ $=\sqrt{p}\begin{bmatrix}\Re\{{\bm{s}}\}\\ \Im\{{\bm{s}}\}\end{bmatrix}\in\mathbb{R}^{2K},$ (23)
${\tilde{{\bm{H}}}}$ =[$=\begin{bmatrix}\Re\{{\bm{H}}\}&-\Im\{{\bm{H}}\}\\ \Im\{{\bm{H}}\}&\Re\{{\bm{H}}\}\end{bmatrix}\in\mathbb{R}^{2K\times 2K},$ (24)
${\bm{\nu}}$ =[$=\begin{bmatrix}\Re\{\bm{\zeta}\}\\ \Im\{\bm{\zeta}\}\end{bmatrix}\in\mathbb{R}^{2K}.$ (25)
The receiver tries to obtain the information signals, using a decoding functionality $f(\cdot)$, which reverts the mapping carried out at the transmitter, either via conventional digital communication-based methods, or the DNN-based algorithms as in [18]. Accordingly, a noisy version of the ground-truth batch, denoted by $\widehat{\bm{x}}$ would be obtained as
$\widehat{\bm{x}}=f(\eta(g(\bm{x})))\approx\bm{x}_{0}+\bm{z},$ (26)
where $\bm{z}$ denotes the unknown reconstruction error.
Remark 1
Notably, in extreme conditions due to low SNR regimes or highly mismatched transceiver and receiver due to large impairment levels, (where even error correction capabilities might not be satisfactory), the idea is to leverage the “denoising-and-generation” characteristics of DDPMs to help facilitate obtaining the ground-truth information signals.
Remark 2
The vanilla format of denoising process of DDPMs, as introduced in Section II, assumes starting the denoising-and-generation process with an isotropic Gaussian noise. However, in a practical wireless system, the reconstructed signals at the receiver are a degraded version, $\widehat{\bm{x}}$, of the ground-truth information signals, not necessarily a pure isotropic Gaussian noise. This degradation is different in different distortion conditions such as different SNRs and impairment levels.
To address the abovementioned challenges, we incorporate the noisy degraded signals into the denoising framework of the diffusion model, realizing a conditional DDPM. Accordingly, the denoising process is modified and the corresponding conditional diffusion and reverse processes are derived.
III-B Proposed Solution: Conditional DDPM for Enhanced Data Reconstruction
Figure 2: Schematic of the Conditional DDPM Framework.
III-B1 Conditional Forward Process
The idea of the employed conditional DDPM is to first incorporate the noisy degraded samples into the forward diffusion framework, so that we can learn, in the reverse diffusion, to remove the exact noise from $\widehat{\bm{x}}$ and refine the ground-truth information signals. For this, a weighted sum of the clean version of ground-truth samples ${\bm{x}}_{0}$ and the noisy samples $\widehat{\bm{x}}$ are considered, using an interpolation hyperparameter $\lambda_{t}$, as shown in Fig. 2. Then the Markov chain-like transition model $q(x_{t}|x_{t-1})$ in (1) is modified into a conditional forward diffusion process as follows.
$q_{\sf{cond}}({\bm{x}}_{t}|{\bm{x}}_{0},{\widehat{\bm{x}}})\sim\mathcal{N}\Big{(}{\bm{x}}_{t};(1-\lambda_{t})\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}+\lambda_{t}\sqrt{\bar{\alpha}_{t}}\widehat{\bm{x}},\delta_{t}I\Big{)},$ (27)
Invoking (27), one can observe that this is a generalization of the vanilla forward diffusion process in (5), as we assume that the mean vector reflects an interpolation between the clean information signals ${\bm{x}}_{0}$ and their noisy degraded version $\widehat{\bm{x}}$. Accordingly, $\lambda_{t}$ is set to start from 00 and gradually increased to $\lambda_{T}\approx 1$. Intuitively, this guides the diffusion model to pay more attention to learning the denoising process from noisy samples over time. According to (27), one can marginalize the conditional distributions over $\widehat{\bm{x}}$ and write
[…]
In order for the conditional diffusion $q_{\sf{cond}}({\bm{x}}_{t}|{\bm{x}}_{0})$ to become a generalization of the vanilla diffusion process $q({\bm{x}}_{t}|{\bm{x}}_{0})$ in (5), one needs to set
$\delta_{t}:=(1-\bar{\alpha}_{t})-\lambda_{t}^{2}\bar{\alpha}_{t}.$ (29)
III-C Conditional Reverse Process
The reverse process starts from ${\bm{x}}_{T}$ as the degraded information signal $\widehat{\bm{x}}$ according to (27) while setting $\lambda_{T}=1$, where we have $p_{\sf{cond}}({\bm{x}}_{T}|\widehat{\bm{x}})\sim\mathcal{N}({\bm{x}}_{T},\sqrt{\bar{\alpha}_{T}}\widehat{\bm{x}},\delta_{T}{\bm{I}}).$ Considering a similar approach to (7), the conditional reverse process is considered to predict ${\bm{x}}_{t-1}$ based on ${\bm{x}}_{t}$ and $\widehat{\bm{x}}$. This results in the following conditional reverse transition probability distribution.
$p_{\sf{cond}}({\bm{x}}_{t-1}|{\bm{x}}_{t},\widehat{\bm{x}})\sim\mathcal{N}({\bm{x}}_{t-1};{\bm{\mu}}_{\theta}({\bm{x}}_{t},\widehat{\bm{x}},t),{\delta}_{t}{\bm{I}}),$ (30)
with ${\bm{\mu}}_{\theta}({\bm{x}}_{t},\widehat{\bm{x}},t)$ formulating the estimated mean of the conditional reverse process. The expression for the conditional mean of the reverse process in (30) should be a generalization to the vanilla format, by incorporating the noisy signals $\widehat{\bm{x}}$ into the formulation as well. Therefore, with a similar expression to (14), the mean vector for each step of the conditional denoising can be expressed as a weighted sum of the denoised signal at step $t$, ${\bm{x}}_{t}$, the conditional information regarding the degraded ground-truth, $\widehat{\bm{x}}$, and the estimated noise $\bm{\epsilon}_{\theta}(\cdot,t)$. Hence, we have
${\bm{\mu}}_{\theta}({\bm{x}}_{t},\widehat{\bm{x}},t)=\psi_{x}{\bm{x}}_{t}+\psi_{\hat{x}}{\widehat{\bm{x}}}-\psi_{\epsilon}{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},\widehat{\bm{x}},t),$ (31)
It is shown in [22] that the values of coefficients $\psi_{x},\psi_{\hat{x}},$ and $\psi_{\epsilon}$ can be estimated by solving the evidence lower bound (ELBO) optimization criterion, where we have
$\psi_{x}$ $=\frac{\delta_{t-1}(1-\lambda_{t})}{\delta_{t}(1-\lambda_{t-1})}\sqrt{\alpha}_{t}+(1-\lambda_{t-1})\frac{\delta_{t|t-1}}{\delta_{t}\sqrt{\alpha}_{t}},$ (32)
$\psi_{\hat{x}}$ $=(\lambda_{t-1}\delta_{t}-\frac{\lambda_{t}(1-\lambda_{t})}{1-\lambda_{t-1}}\alpha_{t}\delta_{t-1})\frac{\sqrt{\bar{\alpha}_{t-1}}}{\delta_{t}},$ (33)
$\psi_{\epsilon}$ $=(1-\lambda_{t-1})\frac{\delta_{t|t-1}\sqrt{1-\bar{\alpha}_{t}}}{\delta_{t}\sqrt{\alpha_{t}}},$ (34)
where $\delta_{t|t-1}\overset{\Delta}{=}\delta_{t}-\left(\frac{1-\lambda_{t}}{1-\lambda_{t-1}}\right)^{2}\alpha_{t}\delta_{t-1}$. As such, the loss function for the conditional model is expressed as
$\mathcal{L}_{t}^{\sf cond}=$
$\mathbb{E}_{{\bm{x}}_{0},{\bm{\epsilon}},\widehat{\bm{x}}}\parallel(\frac{\lambda_{t}\sqrt{\bar{\alpha}_{t}}}{\sqrt{1-\bar{\alpha}_{t}}}{({\widehat{\bm{x}}}-{\bm{x}}_{0})}+\frac{\sqrt{\delta_{t}}}{\sqrt{1-\bar{\alpha}_{t}}}{\bm{\epsilon}})-{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},\widehat{\bm{x}},t)\parallel^{2}_{2},$ (35)
wehre $\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})$. Intuitively speaking, the loss function in (35) is designed such that the conditional DDPM learns, over the denoising steps, to estimate both the Gaussian noise $\bm{\epsilon}$ and the residual errors of estimating $\widehat{\bm{x}}$, i.e., ${\widehat{\bm{x}}}-{\bm{x}}_{0}$. Therefore, the the coefficients for $({\widehat{\bm{x}}}-{\bm{x}}_{0})$ and $\bm{\epsilon}$ are chosen similar to the coefficient of ${\widehat{\bm{x}}}$ and the standard deviation of the diffusion process as in (27), respectively. According to (29)–(35), the training and sampling algorithms are summarized in Algorithm 1 and Algorithm 2, respectively. A visual description of the training phase is also exhibited in Fig. 3.
Hyper-parameters: Number of diffusion steps $T$, neural architecture $\bm{\epsilon}_{\theta}(\cdot,\cdot,t)$, variance schedule $\beta_{t}$ and $\bar{\alpha}_{t},$ and conditioning weights $\lambda_{t}$, $\forall t\in[T]$.
Input: Training samples: The pair of ground-truth information signal and its noisy decoded version from a dataset $\cal D$: $(\bm{x}_{0},{\widehat{\bm{x}}})\in\mathcal{D}$
Output: Trained model $\bm{\epsilon}_{\theta}(\cdot,\cdot,t)$.
Algorithm 1 Training algorithm
1:  while the stopping criteria are not met do
2:     Randomly sample $(\bm{x}_{0},{\widehat{\bm{x}}})$ from $\cal D$
3:     Randomly sample $t\sim\mathsf{Unif}[T]$
4:     Randomly sample $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$
5:     Set $\bm{x}_{t}=(1-\lambda_{t})\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}+\lambda_{t}\sqrt{\bar{\alpha}_{t}}\widehat{\bm{x}}+\delta_{t}\bm{\epsilon}$
6:     Take gradient descent step on
7:     $\quad{\nabla}_{\bm{\theta}}\left\|\frac{1}{\sqrt{1-\bar{\alpha}_{t}}}({\lambda_{t}\sqrt{\bar{\alpha}_{t}}}{({\widehat{\bm{x}}}-{\bm{x}}_{0})}+{\sqrt{\delta_{t}}}{\bm{\epsilon}})-{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},\widehat{\bm{x}},t)\right\|^{2}$
8:  end while
Figure 3: General description of training the diffusion model.
Algorithm 2 Sampling algorithm
1:  Sample $\bm{x}_{T}\sim\mathcal{N}(\bm{x}_{T};\sqrt{\bar{\alpha}_{T}}\widehat{\bm{x}},\delta_{T}\bm{I})$
2:  for $t=T,...,1$ do
3:     $\bm{z}\sim\mathcal{N}(\bm{0},\bm{I})$
4:     $\bm{x}_{t-1}=\psi_{x}{\bm{x}}_{t}+\psi_{\hat{x}}{\widehat{\bm{x}}}-\psi_{\epsilon}{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},\widehat{\bm{x}},t)+\sqrt{\delta_{t}}\bm{z}$
5:  end for
6:  return $\bm{x}_{0}$
(a) Original images $\bm{x}_{0}$
(b) Degraded images $\widehat{\bm{x}}$
(c) Conditional DDPM’s output
Figure 4: Data visualization for the performance evaluation of the proposed scheme for image reconstruction enhancement. From left to right, the ground-truth information signals $\bm{x}_{0}$, the degraded version $\widehat{\bm{x}}$ of the signals decoded at the output of the Sionna’s simulator (before DDPM-based reconstruction), and the final reconstructed signals via our method are shown, respectively.
(a) DNN-based receiver’s output
(b) Conditional DDPM-aided receiver’s output
Figure 5: Comparison between the DNN-based receiver [18] and our DDPM-aided model for image reconstruction.
IV Numerical Results
In this section, we provide different numerical evaluations to highlight the performance of our proposed approach in different scenarios and compared to different baselines. We also evaluate the resilience of our proposed scheme under low-SNR regimes. Our scheme is implemented in conjunction with the NVIDIA Sionna simulator [20] to show the compatibility of the scheme with a practical communication system. To elaborate, after that the diffusion model is trained offline using Algorithm 1, it is plugged into the end of the receiver block of the Sionna simulator, where both scenarios of DNN-based receiver, known as Sionna’s neural receiver, and the conventional receiver are considered in the following evaluations. [^3] One interesting direction would be to propose a framework in which the forward and reverse process of diffusion training are directly incorporated into the communication simulator. This requires modeling the conditional probability distribution of data transmission and reconstruction as modified forward and reverse diffusion processes, respectively, which will be studied in our future works. We consider 64-QAM modulation symbols with 5G low density parity check coding (LDPC) [20] for the modulation and coding of the transmission scheme. We also note that the transmit SNR is defined as
$\Gamma=10\log_{10}(\frac{P}{\sigma^{2}})\hskip 5.69054pt\text{dB}.$ (36)
Without loss of generality, we normalize the transmit power by setting the average signal power to $P=1$ for all experiments, while varying the SNR by setting the standard deviation (std) of noise $\sigma$ [25].
For the choice of neural network for our conditional diffusion model, we follow the model employed in [7], and adopt the U-Net architecture [23] as our denoising neural network ${\bm{\epsilon}}_{\theta}({\bm{x}}_{t},\widehat{\bm{x}},t)$ with input size 32 and depth of 4. The employed neural network is supposed to take the distorted image at each time step, and output the predicted noise with the same size as the input image. The U-Net architecture matches well with this requirement. It consists of a contracting path and an expansive path, which respectively capture contextual information and reduce the spatial resolution of the input, and then upsamples the feature maps, while also performing convolutional operations [23]. Additionally, it introduced residual connections between the encoder and decoder, which greatly improves the gradient flow. Evaluations are carried out on MNIST dataset (resized to $32\times 32$) which contains 60,000 training images and 10,000 testing images of hand-written digits. The corresponding dataset $\mathcal{D}$ in Algorithm 1, is then obtained by transmitting the training samples of MNIST over the Sionna’s simulator and obtain the decoded signals at the receiver’s output, and then pairing each of the ground-truth samples with their degraded version to create $(\bm{x}_{0},{\widehat{\bm{x}}})\in\mathcal{D}$. The training is carried out for 10 epochs over $T=200$ steps, and using adaptive moment estimation (Adam) optimizer with learning rate $\lambda=10^{-3}$. To stabilize the training process, exponential moving average (EMA) method is implemented [24] for model update during training. This helps maintain a form of “model momentum.” Specifically, instead of directly updating model’s weights, a copy of the previous weights is kept, and the weighted average between the previous and new version of the weights are calculated for the model update. We also set $T=200$. For the conditioning weights $10<\lambda_{t}<$ 1, we consider a linear scheduling over $T=200$ steps. For the variance scheduling $\beta_{t}$, we also consider a linear scheduling, starting from $\beta_{0}=0.0001$ to $\beta_{T}=0.0095$. We note that the values for the variance scheduling were obtained based on trial-and-error over different scheduling formats. To the best of our knowledge, there does not exist a unified generic solution for the choices of variance scheduling in the diffusion model literature, and it depends on the specific problem to which the diffusion model is applied. To enable employing only one neural network for the entire denoising steps, instead of training $T$ distinct models for each time-step $t\in[T]$, we share the parameters of the neural network across time-steps. More specifically, as shown in Fig. 3, we encode the time-steps $t\in[T]$ and input it to each hidden layer of our neural network as a positional embedding [7]. The output of the hidden layers are then multiplied by the time embeddings [7, 9, 12]. Intuitively, this makes the model “know” at which particular time-step it is operating for every sample in the batch.
IV-A Data Visualization
Fig. 4 visualizes the performance of our conditional DDPM-aided reconstruction enhancement scheme. For this figure, we consider the extreme condition of $\Gamma=0$ dB channel SNR, with the transmitter’s impairment level $\kappa^{t}=0.1$ and the receiver’s impairment level of $\kappa^{r}=0.15$. As can be seen from the figure, under such extreme conditions, the naive format of a communication system cannot perform well in removing the noise and distortions to obtain a clean denoised version of the ground-truth information signals (Fig. 4-(b)). In such scenarios, our proposed diffusion-aided scheme can improve the reconstruction of images by removing the noise and distortions conditioned on the degraded signals. This results in a significant improvement in the decoding performance. This improvement is actually obtained via exploiting the prior knowledge learnt by our conditional DDPM. One might argue that in such cases, we can play with the coding rate and employ stronger error correction codes with lower code rates. However, this can significantly reduce the data rate, while stronger error correction codes might not necessarily work well under such extreme conditions either. We note that in the subsequent figures, we numerically study this, and we show that the DDPM outperforms those schemes in low SNR regimes.
Fig. 5 visualizes the output of our proposed diffusion-aided scheme, compared to a DNN-based receiver as benchmark [18]. For this experiment, we set $\Gamma=5$ dB SNR and the transmitter’s impairment level $\kappa^{t}=0.1$ and the receiver’s impairment level of $\kappa^{r}=0.15$. Notably, compared to the DNN benchmark our scheme highlights a considerably better performance in obtaining clean samples of the information signals, by removing noise and other residual decoding errors. This is due to the fact that the diffusion-based method exploits the generative prior knowledge of the underlying distribution. However, a DNN-based receiver as proposed in [18] does not take the “content”, nor any prior knowledge of the information signals into account when decoding the information bits. We further emphasize that numerical comparisons between our conditional DDPM and the DNN benchmark are provided in the subsequent figures.
IV-B Reconstruction Performance Metrics
In this subsection, we numerically assess the reconstruction performance of the proposed scheme in different scenarios.
Figure 6: Reconstruction performance in terms of peak SNR (PSNR) for different training strategies.
Fig. 6 shows the reconstruction performance in terms of peak signal-to-noise ratio (PSNR) metric for our proposed scheme. For case I to III, we consider $(\kappa^{t},\kappa^{r})=(0.01,0.01)$, $(\kappa^{t},\kappa^{r})=(0.05,0.1)$, and $(\kappa^{t},\kappa^{r})=(0.05,0.15)$, respectively. In this figure, we study the inference performance of the employed conditional DDPM when trained with two different training strategies: varying training SNR, or fixed training SNR. As mentioned earlier, the training samples are obtained by transmitting the training set of MNIST over the Sionna’s simulator (treated as the samples of ${\bm{x}}_{0}$) and obtain the decoded signals at the receiver’s output (treated as the samples of $\widehat{\bm{x}}_{0}$). This can be done with either fixing the communication scenario to a determined SNR, or simulating the system with different transmit SNRs. For the varying scenario, we set $-15\leq\Gamma^{\sf train}\leq 15$ dB by trail and error over different ranges of SNRs. Notably, when the conditional diffusion model is trained using training samples that are obtained over a wide range of communication SNRs, this makes the system more robust to the unseen scenarios in the inference phase. Therefore, as can be seen from the figure, higher PSNR values can be achieved in this case, while the performance of the fixed training SNR scenario saturates too early. We further note that this saturation is also due to another fact as well, which is known as “ceiling phenomena” in the literature of hardware-impaired communications [21]: By increasing the SNR, the effect of channel noise becomes negligible compared to the effect of hardware impairments, hence the performance metrics become saturated. As can be seen from the figure, the PSNR increases with the increase in SNR, due to having less noise communication, which makes it easier for the employed diffusion to denoise and reconstruct when facing with a less degraded signal. In addition, lower impairment levels indicate higher qualities for the transceiver hardware which results in less mismatches between the transmitted signals and the decoded ones, thus improving the reconstruction performance. Generally speaking, the proposed scheme shows a near-invariant performance with respect to the hardware impairments (as we can see the slight change in PSNR with respect to the three scenarios). This is further studied in the subsequent figures as well.
Figure 7: Reconstruction performance in terms of peak SNR (PSNR): Comparison between generative (DDPM-based) learning approach vs. discriminative (DNN-based) learning approach.
Figure 8: Reconstruction performance in terms of MSE between the original images and the reconstructed ones.
Fig. 7 evaluates the reconstruction performance of our proposed approach when it is employed in conjunction with the Sionna’s neural receiver [^4] https://nvlabs.github.io/sionna/examples/Sionna_tutorial_part4.html, compared to the non-DDPM approach which simply employs the Sionna’s neural receiver without any diffusion enhancement. For the DNN-based receiver, we consider the network architecture proposed in [18], trained for 5000 iterations with learning rate $0.01$. Since we have employed our diffusion model at the receiver, we only exploit the receiver DNN of [18], and fine-tune it for benchmarking, so that we can have a fair comparison. For the DNN benchmark, three hidden linear layers with $64$ neurons and rectified linear unit (ReLU) activation functions are considered. Based on our trail and errors, increasing the depth of the conventional DNN, or the number of hidden neurons does not result in any significant improvement in its performance. The figure highlights about $10$ dB improvement in low-SNR regimes, as well as more than $5$ dB improvement in higher SNRs, when using our generative model. This is due to the fact that our DDPM-based scheme takes into account the entire underlying distribution of the ground-truth signals that are supposed to be sent over the wireless channel. However, the conventional DNNs neglect the underlying “content” of the information signals, and simply care about the “bit-wise” transmission, using discriminative loss functions, such as binary cross-entropy, rather than a content-wise” approach.
Fig. 8 demonstrates the image reconstruction error, in terms of the mean squared error (MSE) versus SNR for $\kappa^{t}=\kappa^{r}=0.25$. The figure shows that under extreme conditions (low-SNR and mid-SNR regimes and highly-impaired communications), a naive scheme (without our proposed diffusion model) cannot perform well in terms of the image reconstruction, even if the error correction rate is decreased from $0.5$ to $0.2$. When the conditional DDPM is employed, the reconstruction is enhanced significantly, by more than 10 dB. This is due to the fact that our scheme leverages the denoising capabilities of diffusion models, as well as being conditioned on the degraded output of a naive communication system, which guides the diffusion to improve the data decoding and reconstruction as shown in the figure. The figure suggests that instead of employing error correction codes with lower code rates, to improve the data decoding quality, our proposed scheme can be employed while preserving the data rate. As can be seen from the figure, more than $5$ dB performance improvement, as well as $2.5$ times higher data rates can be achieved under low SNR regimes with our proposed scheme, instead of reducing the code rate from $0.5$ to $0.2$ in a naive scheme. This is applicable to scenarios in which data rates and reliable communications are assumed to be taken into account at the same time. Also note that although the gap between a naive scheme with lower code rate (blue curve) and the DDPM-aided scheme with code rate of $0.5$ (black curve) decreases as the SNR increases, the figure shows that if we employ the DDPM, the same performance can be achieved, while we do not even need to decrease the data rate.
Figure 9: Reconstruction performance between the original images and the reconstructed ones for different impairment levels: Diffusion models vs. channel codes.
Fig. 9 shows the reconstruction performance between the original images and the reconstructed ones for two different impairment levels. In this experiment, we aim to show that the functionality of the proposed conditional diffusion model can be exploited to compensate for the underlying errors within the noisy reconstructed images obtained at the receiver, instead of employing channel codes. As can be seen from the figure, by employing our conditional DDPM, the quality of the information signals (images in our scenario) can be improved more than the case where error correction codes are employed. For instance, by employing our proposed scheme, more than $10$ dB improvement in reconstruction performance can be achieved at $5$ dB SNR and impairment level $0.1$. This can be explained due to the fact that channel codes do not take into account the underlying structure and the contents of the information signals that are transmitted over the channel. However, our proposed scheme essentially considers the content of the transmitted signals when carrying out the denoising and reconstruction. Also note that error correction codes require sending redundancy bits which decrease the data rate, while our diffusion-based scheme has achieved better performance without the need for transmitting such redundancy bits. Therefore, the figure indicates that our scheme can improve the reliability of the communication systems, while saving the resources for only transmitting information bits. Obviously, this performance improvement comes with a trade-off in the sense that the DDPM model requires higher computation resources than the error correction codes. Nevertheless, one should note that the computations related to error correction codes are implemented in both of the communication sides, i.e., the transmitter (the encoder part), and the receiver (the decoder part), while the proposed diffusion is only employed and run at the receiver side. We note that studying the mathematical models and expressions for this trade-off is left for our future works.
Figure 10: MSE between the original images and the reconstructed ones for different levels of hardware impairment.
Fig. 10 studies the effect of different HWI levels on the reconstruction performance of our scheme over Rayleigh fading channels. For this figure, we set $\kappa^{t}=0.05$ and vary the impairment level of the receiver, $\kappa^{r}$, over the typical ranges specified in Ref. [21]. The reconstruction results are obtained in terms of MSE metric over $20$ realizations of the system. The figure highlights an important characteristic of our proposed scheme. Our DDPM-based communication system is resilient against hardware distortions, since the reconstruction performance does not change with the increase in the impairment level. Notably, our scheme showcases a near-invariant reconstruction performance with respect to noise and impairment levels. This is achieved due to the so-called “variance scheduling” of diffusion framework in (3), which allows the system to become robust against a wide range of distortions caused by channel and hardware impairments.
V Conclusions & Future Works
In this paper, we have proposed conditional DDPMs to enhance the data reconstruction in wireless communication schemes. The key idea is to leverage the generative prior of diffusion models to learn a noisy-to-clean transformation of the information signal, conditioned on the degraded version of the receiver data. The proposed scheme is applicable to scenarios in which a prior knowledge of the information signals is available, e.g., in multimedia transmission. Our numerical results on MNIST dataset have shown that instead of employing complicated channel codes that reduce the information rate, conditional diffusions can be employed that improve reliable data transmission, especially under extreme channel conditions due to low signal-to-noise ratio (SNR), or hardware-impaired communications. Our evaluations have been carried out using the NVIDIA Sionna simulator, where it has been shown that more than $10$ dB improvement as well as $2$ times higher data rates can be achieved under low SNR regimes with our proposed scheme, compared to a conventional error correction-based data transmission and reconstruction. Moreover, we highlight the outperformance of the proposed generative learning approach compared to the conventional contrastive learning models based on DNNs. One important direction to be studied in the future works is to analytically study the complexity of generative AI-based approach versus the normal heavily channel coded approach.